# Heterogeneous Graph Neural Networks for Data-driven Traffic Assignment

### v1

@@ Abstract
The traffic assignment problem is one of the significant components of traffic flow analysis for which various solution approaches have been proposed. However, deploying these approaches for large-scale networks poses significant challenges. In this paper, we leverage the power of heterogeneous graph neural networks to propose a novel data-driven approach for traffic assignment and traffic flow learning. The proposed model is capable of capturing spatial traffic patterns across different links, yielding highly accurate results. We present numerical experiments on urban transportation networks and show that the proposed heterogeneous graph neural network model outperforms other conventional neural network models in terms of convergence rate, training loss, and prediction accuracy. Notably, the proposed heterogeneous graph neural network model can also be generalized to different network topologies. This approach offers a promising solution for complex traffic flow analysis and prediction, enhancing our understanding and management of a wide range of transportation systems.
@@
traffic assignment, graph neural network, traffic flow prediction, flow conservation
@@ 1 Introduction
The traffic assignment problem, as a significant component of transportation network performance analysis, helps understand spatial and temporal traffic flow patterns for effective transportation network management and provides insights into bottlenecks, and traffic congestion (Nie et al., 2004). Congestion is a major concern in urban areas, leading to increased travel times, fuel consumption, and environmental pollution. By identifying the bottlenecks in transportation networks, city planners can make a strategic plan to increase road capacities to mitigate traffic congestion.
To this end, the objective of the traffic assignment problem is to determine the traffic flow distribution and identify traffic bottlenecks on a given road network. Numerous mathematical models have been proposed to tackle traffic assignment problems. The user equilibrium (UE) principle and system optimum (SO) principles are two fundamental but effective solutions to traffic assignment problems. The UE principle, also known as Wardrop’s first principle, states that drivers between each origin-destination (OD) pair cannot unilaterally reduce travel costs by unilaterally shifting to another route (Kuang et al., 2021). The system optimum assignment, which is based on Wardrop’s second principle, emphasizes driver cooperation to minimize the total system travel time (Seliverstov et al., 2017). These models have been used extensively in practice to estimate traffic flow patterns and optimize transportation infrastructure investments. However, these approaches are computationally expensive for a large-scale network. Substantial efforts have been made to find efficient algorithms to solve this problem (Nie et al., 2004; Di Lorenzo et al., 2015). Furthermore, researchers have proposed various extensions to the UE and SO models, such as the stochastic user equilibrium (Damberg et al., 1996), multi-modal traffic assignment (Pi et al., 2019), and dynamic traffic assignment (Ben-Akiva et al., 2012). These models incorporate additional factors, such as travel time variability, driver behaviors, and network capacity constraints, to provide more realistic traffic assignment solutions. However, there is still a pressing need for more effective and accurate methods for traffic assignment problems, particularly for traffic planning and optimization.
To perform transportation performance analysis, OD demand needs first to be obtained, which is a significant component in traffic assignment problems. OD demand estimation has been widely studied in the transportation engineering literature. The statistical model is one of the common approaches for OD estimation. It leverages multiple data sources to estimate vehicle flow between different locations in a network (Hazelton, 2008; Jin et al., 2014). Furthermore, the increasing availability of crowdsourced data, including GPS data and social media, enables us to extract OD demand and traffic flow data at a large scale and over extended periods (Toole et al., 2015; Zhang et al., 2020). Besides, neural network approaches (Tang et al., 2021; Zhang et al., 2021) have also been applied to estimate OD estimation, which allows for the extraction of complex patterns from high-dimensional and noisy transportation data.
In recent years, data-driven approaches have leveraged the availability of large-scale transportation data and computational resources in various transportation problems. For instance, the recurrent neural network (RNN), including the long short-term memory (LSTM) and gated recurrent unit (GRU) networks, have been applied to different prediction tasks in transportation (Zhaowei et al., 2020; Fang et al., 2022). However, standalone RNNs lack the ability to learn spatial correlation. To overcome this weakness and capture the spatio-temporal relationship of traffic flows at different locations, Ye et al. (2019) combines convolutional neural networks (CNN) with LSTM to predict multiple transportation demands using real-world taxi and sharing bike demand data. Recently, the graph neural network (GNN) has emerged as a powerful tool in transportation performance analysis because of its capability to effectively model and analyze data represented in the form of a graph. Compared with RNNs and CNNs, a key advantage of GNNs is their ability to capture spatial and relational information inherent in transportation networks.
Graph neural networks have been shown to be effective in network modeling and management of transportation systems (Rahman and Hasan, 2022; Liu and Meidani, 2023a). However, these models encountered a few limitations. Firstly, the proposed GNN models can only be trained on a fixed topology, and need to be re-trained whenever the topology of the road network changes. Additionally, these studies didn’t adequately consider the dynamic nature of traffic flow, e.g., when link capacities alter due to traffic accidents or damages to roadways. These drawbacks limit the adaptability and applicability of these models in real-world scenarios.
To address this challenge, in this paper, we propose a novel GNN-based approach for static traffic assignment. Specifically, we focus on predicting the flow-capacity ratio of different scenarios in the context of traffic assignment. In particular, to address the long-range effects caused by OD pairs that are multiple-hops apart, we create a heterogeneous GNN that consists of additional "virtual" links. The major contributions of this work are as follows: (1) a heterogeneous GNN model is proposed with "virtual" OD-links and the "real" roadway links , which incorporates the attention-based mechanism and message-passing mechanism; (2) the proposed model also integrates a physics-based loss based on the flow conservation law into the total loss function to accelerate the learning process; (3) the proposed model is generalizable to different road network topologies, link characteristics, and OD demands.
@@
The remainder of this article is structured as follows. General backgrounds on the traffic assignment problem, the neural network and graph neural network models are presented in Section 2. Section 3 includes the explanation of the proposed heterogeneous graph neural network for traffic flow learning. Furthermore, the experiments with urban road networks and generalized synthetic networks are presented to demonstrate the accuracy and generalization capability of the proposed framework in Section 4. Finally, the conclusion and discussion of the proposed framework are presented in Section 5.
@@ 2.1 Traffic Assignment Problem
where $f_{e}$ and $Z_{e}(f_{e})$ are the total flow and the link cost function on link $e$, respectively. The link cost function can be expressed in terms of travel time, travel distance, or other relevant factors. Therefore, the objective function is to minimize the summation of link cost over the graph. The traffic assignment problem can have different formulations depending on the specific objectives and constraints considered in the optimization process. For instance, under user equilibrium formulation, the traffic assignment problem could be solved by optimizing Beckmann’s formulation (Beckmann et al., 1956):
$\min:\quad$ $z(x)=\sum_{e\in\mathcal{E}}\int_{0}^{x_{e}}t_{e}(\omega)\mathrm{d}\omega$ (2)
$\mathrm{s.t.}\quad$ $\sum_{k}f_{k}^{rs}=q_{rs},\ \forall r,s\in\mathcal{V},$
@@
where $t_{e}(f_{e})$ is the link travel time function, $q_{rs}$ is the total demand from source $r$ to destination $s$, $f^{rs}_{k}$ represents the flow on $k^{\mathrm{th}}$ path from $r$ to s. $\zeta^{rs}_{e,k}$ is the binary value, which equals to 1 when link $e$ is on $k^{\mathrm{th}}$ connecting $r$ and $s$.
2.2 Neural Network
@@ 2.3 Graph Neural Network
3 Traffic Flow Learning using Heterogeneous Graph Neural Networks
In this section, we elaborate on the proposed architecture of the heterogeneous graph neural network for traffic flow learning. The detail of the proposed model is shown in Fig. 1. It consists of the embedding preprocessing block, the heterogeneous graph attention block, and the prediction block. The details of each module are described as follows.
Figure 1: The illustration of the heterogeneous graph neural network for traffic flow learning. The original node feature is transformed using multi-layer perception. Then, the node feature and edge feature are first passed through the virtual links and then passed through the real links. The flow-capacity ratio of each link is calculated using the source node feature, destination node feature, and edge feature.
@@ 3.1 Graph Construction & Feature Preprocessing
The heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{r},\mathcal{E}_{v})$ for traffic flow learning consists of one type of node representing the intersections of road segments and two edge types: real links and virtual links. The real links represent the road segments in the road network, while the virtual links represent the auxiliary link between the origin and destination nodes. The auxiliary links can be considered as the edge augmentation technique to enhance the feature updating. The node features $\bm{X}_{n}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|+2}$ are the origin-destination demand of each node and the node geocoordinates. The real link features $\bm{X}_{e,r}\in\mathbb{R}^{|\mathcal{E}_{r}|\times 2}$ include the free-flow travel time and the link capacity, while the virtual link features $\bm{X}_{e,v}\in\mathbb{R}^{|\mathcal{E}_{v}|\times 1}$ are binary values indicating the existence of OD demand between origin and destination nodes.
Furthermore, the original node features are often sparse and non-normalized. To address this issue, we employ a preprocessing step to encode the raw features into a lower-dimensional representation that captures the essential characteristics of the data while preserving the semantic information. The generated node feature embedding size is $\bm{X}^{0}_{n}\in\mathbb{R}^{|\mathcal{V}|\times N_{v}}$, where $N_{v}$ is the embedding size. Similarly, the edge features are also normalized before being propagated in the message passing. The normalized edge feature of real edges and virtual edges are denoted as $\bm{X}^{0}_{e,r}\in\mathbb{R}^{|\mathcal{E}_{r}|\times 2}$ and $\bm{X}^{0}_{e,v}\in\mathbb{R}^{|\mathcal{E}_{r}|\times 1}$, respectively. Additionally, it should be noted that there is an overlap between the real links and virtual links in the heterogeneous graph because it is possible that a road and an OD demand both exist between the same node pair.
@@ 3.2 Graph Spatial Features Extraction
As discussed in Section 2.3, effectively aggregating information from nodes and edges of different types is a major challenge for heterogeneous graphs. In order to address this challenge, we propose a novel approach that leverages attention mechanisms on both real links and virtual links. This allows our model to selectively focus on the most relevant information from each edge type, improving the overall performance of the model.
As the first step of spatial feature extraction, to enhance the modeling capability and capture more fine-grained relationships, we employ a key-query-value mechanism on node features in addition to the weighted sum operation. For single-head attention, the node features are transformed into key, query, and value matrices through linear transformations:
$\bm{K},\bm{Q},\bm{V}=[\bm{W}_{Q}\bm{X}^{0}_{n},\bm{W}_{K}\bm{X}^{0}_{n},\bm{W}_{V}\bm{X}^{0}_{n}],$ (7)
where $\bm{W}_{K}\in\mathbb{R}^{N_{v}\times N_{h}}$, $\bm{W}_{Q}\in\mathbb{R}^{N_{v}\times N_{h}}$, and $\bm{W}_{V}\in\mathbb{R}^{N_{v}\times N_{h}}$ are learnable weight matrices. $N_{h}$ is the output size of the linear transformation. The value matrix captures the learned representations and importance scores for each node. The query matrix represents the target node’s characteristics, while the key matrix contains the features of neighboring nodes. The attention scores between the query vector and each key are calculated with the dot product, normalized by a softmax function to ensure that the weights sum up to one:
$s_{ji}=\frac{\exp(\bm{q}_{j}^{T}\bm{k}_{i})}{\sum_{k\in\mathcal{N}(i)}\exp(\bm{q}_{k}^{T}\bm{k}_{i})},$ (8)
where $\bm{k}_{i}$ and $\bm{q}_{j}$ denote the $i^{\mathrm{th}}$ and $j^{\mathrm{th}}$ column of matrix $\bm{K}$ and $\bm{Q}$, respectively. The attention scores are then used in the weighted sum of the corresponding values, allowing the model to focus on more informative features. When the edge consists of multiple features, the output of the attention layer $\bm{R}\in\mathbb{R}^{|\mathcal{V}|\times(N_{h}\times h)}$ is calculated using the attention score from Equation 8, value vector, and edge features from the neighbor of node $i$:
$\bm{R}=\bigoplus_{i=1}^{N}\left(\sum_{l=1}^{L}\sum_{k\in\mathcal{N}(i)}s_{ik}\bm{v}_{k}x_{ik,l}\right),$ (9)
where $\bm{v}_{k}$ denote the $k^{\mathrm{th}}$ column of value matrix $\bm{V}$, $x_{ik,l}$ represent the $l^{\mathrm{th}}$ edge feature on link $(i,k)$. This allows the model to focus on the most relevant output vectors for a given input vector. For multi-head attention, the process of Equation 7, 8, and 9 is repeated $h$ times to get multiple head attention output $[\bm{R}^{(0)},\bm{R}^{(1)},,\bm{R}^{(h-1)}]$, where $h$ is the number of heads. The learned embedding of original node features is then passed into a linear transformation layer. We concatenate the output tensors $\bm{R}^{(0)},\bm{R}^{(1)},\dots,\bm{R}^{(h-1)}$ along the feature dimension and multiply a learnable weight matrix to produce the final output tensor $\bm{O}$:
$\bm{O}=\bigoplus_{i=0}^{h-1}\left(\bm{W}_{o}\bm{R}^{(i)}\right).$ (10)
where $\bm{W}_{o}\in\mathbb{R}^{(N_{h}\times h)\times N_{o}}$, $N_{o}$ is the output dimension of the final node embedding of a single layer GNN.
As shown in Figure 1, multiple GNN layers are stacked to extract spatial features. The first GNN layer aggregates the node features along the virtual links. While the following GNN layers aggregate the node features along the real link. Compared with the homogeneous graph constructed with only real links, the link augmentation and message passing through virtual links can be considered as a dimension-reduction technique to reduce the number of hops required for distant nodes to gather messages. As a result, it requires fewer GNN layers for effective feature aggregation and following edge prediction in the heterogeneous graph.

### v2

# Heterogeneous Graph Neural Networks for End-to-End Traffic Assignment and Traffic Flow Learning

@@ Abstract
The traffic assignment problem is one of the significant components of traffic flow analysis for which various solution approaches have been proposed. However, deploying these approaches for large-scale networks poses significant challenges. In this paper, we leverage the power of heterogeneous graph neural networks to propose a novel data-driven approach for end-to-end traffic assignment and traffic flow learning. Our model integrates an adaptive graph attention mechanism with auxiliary "virtual" links connecting origin-destination node pairs, This integration enables the model to capture spatial traffic patterns across different links, By incorporating the node-based flow conservation law into the overall loss function, the model ensures the prediction results in compliance with flow conservation principles, resulting in highly accurate predictions for both link flow and flow-capacity ratios. We present numerical experiments on urban transportation networks and show that the proposed heterogeneous graph neural network model outperforms other conventional neural network models in terms of convergence rate and prediction accuracy. Notably, by introducing two different training strategies, the proposed heterogeneous graph neural network model can also be generalized to different network topologies. This approach offers a promising solution for complex traffic flow analysis and prediction, enhancing our understanding and management of a wide range of transportation systems.
@@
traffic assignment problem , graph neural network , traffic flow prediction , flow conservation , heterogeneous graph
journal: Transportation Research Part C: Emerging Technologies
@@ 1 Introduction
Traffic assignment plays a significant role in transportation network analysis. Solving the traffic assignment problem (TAP) enables a deeper understanding of traffic flow patterns and provides insights into traffic congestion (Nie et al., 2004). Traffic congestion is a major concern in urban areas, leading to increased travel times, fuel consumption, and environmental pollution. By identifying the bottlenecks in transportation networks, city planners can make a strategic plan to increase road capacities to mitigate traffic congestion. To this end, the primary objective of the traffic assignment problem is to determine the traffic flow distribution and identify traffic bottlenecks on a given road network. For solving traffic assignment problems, there are two major formulations with different assumptions: (1) the user equilibrium (UE) assignment and (2) the system optimum (SO) assignment. The UE assignment identifies an equilibrium state in which drivers between each origin-destination (OD) pair cannot reduce travel costs by unilaterally shifting to another route (Kuang et al., 2021). On the other hand, in the SO assignment, drivers choose the routes collaboratively so that the total system travel time is minimized (Seliverstov et al., 2017).
Traffic assignment problems have appeared in other transportation engineering problems. For instance, city planners leverage traffic assignment to evaluate the network performance under extreme event scenarios, including hurricanes (Zou and Chen, 2020) and earthquake (Silva-Lopez et al., 2022; Liu and Meidani, 2023a). These evaluations ensure the transportation infrastructure aligns with the city’s evolving needs and requirements. Furthermore, the UE-TAP is employed in road network design and retrofit investment optimization as a lower-level constraint in a bi-level optimization problem (Hu et al., 2021; Madadi and de Almeida Correia, 2024).
As a critical preliminary step in setting up the TAPs, it is essential to first obtain accurate origin-destination demands that reflect the realistic travel demands. Multiple research has been proposed to estimate regional OD demands (Zhang et al., 2020; Tang et al., 2021; Zhang et al., 2021), leveraging data from various sensors. Despite these methodological advancements, practical challenges persist. Specifically, sensor failures and malfunctions can impose a significant loss of relevant and reliable OD demand information (Sun et al., 2022). Consequently, the inaccurate estimation of OD demand will hinder the accurate calculation of traffic flows. This issue underscores a research gap in effectively solving TAPs and estimating traffic flows under inaccurate OD demand scenarios. In order to solve TAPs, various efficient algorithms have been proposed (Nie et al., 2004; Di Lorenzo et al., 2015). Despite these advances, it is still computationally demanding to conduct a robust network planning for large-scale networks, when multiple relevant scenarios are considered. This highlights the need to develop scalable solution approaches for TAPs to accommodate diverse network scenarios. In recent years, data-driven approaches have leveraged the availability of large-scale transportation data in various transportation problems. For instance, Ye et al. (2019) combines convolutional neural networks to predict multiple transportation demands using real-world taxi and shared bike demand data. Furthermore, recurrent neural networks (RNNs) have been applied to different prediction tasks in transportation engineering (Zhaowei et al., 2020; Fang et al., 2022). Also, Liu et al. (2023) proposed an implicit neural network to travel choice preferences and equilibrium state modeling from multi-day link flow observations.
Recently, graph neural networks (GNNs) have emerged as a powerful tool in transportation performance analysis since they propagate the node features through a graph structure. Compared with RNNs and CNNs, a key advantage of GNNs is their ability to capture spatial and relational information inherent in transportation networks. Among the recent works, Nishi et al. (2018) developed a traffic signal control algorithm based on graph convolutional networks and reinforcement learning. Liu and Meidani (2022) utilizes a graph convolutional network to analyze the node-to-node accessibility of regional transportation systems under earthquakes. Despite these advancements in neural networks, the literature on applying neural networks to TAPs remains relatively sparse. Fan et al. (2023) integrated CNNs with RNNs for predicting the traffic flow distribution by considering the transportation network as the grid map. Furthermore, Rahman and Hasan (2023) utilizes the graph convolutional network for solving UE-TAP and predicting the traffic flow pattern. However, the aforementioned model encountered a few limitations. First of all, one major limitation of the grid-based CNN model (Fan et al., 2023) is that it cannot fully capture the topologies of the transportation system. Furthermore, these models don’t adequately consider the transportation network under various scenarios, e.g., link capacity reduction due to traffic accidents or lane closures due to maintenance. Besides, they don’t fully investigate the model performance under the situations of incomplete/missing OD demand information. Moreover, how these models perform under out-of-distribution data is not comprehensively explored. These limitations highlight the need for further research to enhance the adaptability and real-world applicability of GNNs for traffic assignment problems.
It should be noted that in the transportation network, a direct link between the origin and destination nodes in general may not exist. This while the traffic flows do in fact depend on the relationship between the OD pairs and OD demands. This observation motivates us to consider introducing the auxiliary links between the origin and destination nodes within the original transportation network, which could encapsulate valuable supplementary information, which is often neglected in the previous literature. In this paper, we propose a novel heterogeneous GNN model for efficient traffic assignment. In addition to roadway links, our model includes virtual links between origin-destination nodes, enriching the model with comprehensive OD demand information and thereby facilitating enhanced feature propagation across the network. Furthermore, we introduce a novel adaptive graph attention mechanism to propagate the node features efficiently. And, the proposed model transforms node embeddings into link embeddings for link flow and flow-capacity ratio estimations. To summarize, the major contributions of this work are as follows: (1) This is the first GNN learning of UE-TAP that integrates interdependencies between origin and destination nodes via a heterogeneous graph structure that consists of physical and virtual links and an adaptive attention-based mechanism; (2) the proposed model is trained using data and the governing conservation law and as such the estimated flows are more accurate; (3) due to the integration of virtual links and regularization based on the conservation law, the performance on unseen graphs is improved. The efficiency and generalization capability of the proposed model are investigated through multiple experiments with different road network topologies, link characteristics, and OD demands.
@@
The remainder of this article is structured as follows. General backgrounds on the traffic assignment problem, the neural network and graph neural network models are presented in Section 2. Section 3 includes the explanation of the proposed heterogeneous graph neural network for traffic assignment and traffic flow learning. Furthermore, the experiments with urban road networks and generalized synthetic networks are presented to demonstrate the accuracy and generalization capability of the proposed framework in Section 4. Finally, the conclusion and discussion of the proposed framework are presented in Section 5.
@@ 2.1 Traffic Assignment Problem
where $f_{e}$ and $Z_{e}(f_{e})$ are the total flow and the link cost function on link $e$, respectively. The link cost function can be expressed as the function of travel time, travel distance, or other relevant factors. Besides, the traffic assignment problem can have different formulations depending on the specific objectives and assumptions. For instance, the (Beckmann et al., 1956) addresses the user equilibrium traffic assignment problem by optimizing the following objective function:
$\min:$ $z(x)=\sum_{e\in\mathcal{E}}\int_{0}^{x_{e}}t_{e}(\omega)\mathrm{d}\omega$ (2)
$\mathrm{s.t.}$ $\sum_{k}f_{k}^{rs}=q_{rs},\ \forall r,s\in\mathcal{V},$
@@
where the objective function is the summation over all road segments of the integral of the link travel time function between 0 and the link flow. $t_{e}(\cdot)$ is the link travel time function, $q_{rs}$ is the total demand from source $r$ to destination $s$, $f^{rs}_{k}$ represents the flow on $k^{\mathrm{th}}$ path from $r$ to s. $\zeta^{rs}_{e,k}$ is the binary value, which equals 1 when link $e$ is on $k^{\mathrm{th}}$ connecting $r$ and $s$. It is noted that the objective function in Beckmann’s formulation serves more as a mathematical construct for optimization than a direct physical representation. Compared with UE-TAP, SO-TAP changes the objective function to the summation of the travel time of all vehicles, which reflects a system-optimized perspective.
2.2 Neural Networks
@@ 2.3 Graph Neural Network
3 Traffic Assignment and Traffic Flow Learning using Heterogeneous Graph Neural Networks
In this section, we elaborate on the proposed architecture of the heterogeneous graph neural networks for traffic assignment and traffic flow learning. The illustration of the proposed model is shown in Fig. 1. It consists of three modules: graph construction & feature preprocessing module; spatial feature extraction module, and edge prediction module. The detail explanation of each module is described as follows.
Figure 1: The illustration of the heterogeneous graph neural network for traffic flow learning. The proposed model consists of three parts: graph construction & feature preprocessing module; spatial feature extraction module, and edge prediction module. The graph features are first passed into the virtual encoder (V-Encoder) through the virtual links. Then the graph embeddings are passed into the real encoder (R-Encoder) through the real links. The flow-capacity ratio and link flow of each link are calculated using the source node feature, destination node feature, and normalized edge feature.
@@ 3.1 Graph Construction & Feature Preprocessing
The heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{r},\mathcal{E}_{v})$ for traffic flow learning consists of one type of node representing the intersections of road segments and two edge types: real links and virtual links. The real links represent the road segments in the road network, while the virtual links represent the auxiliary link between the origin and destination nodes. The incorporation of the virtual link, as an edge augmentation technique, is strategically employed to facilitate enhanced feature updating. In this model, the node feature attribute is denoted as $\bm{X}_{n}\in\mathbb{R}^{|\mathcal{V}|\times(|\mathcal{V}|+2)}$, where $\mathcal{V}$ represents the graph’s node set. Specifically, each row $\bm{x}_{u}\in\mathbb{R}^{1\times(|\mathcal{V}|+2)}$ within this matrix corresponds to the feature vector of a single node $u\in\mathcal{V}$ including the origin-destination demand as well as geographical coordinates. The feature representation for the entire set of real edges is represented with $\bm{X}_{e,r}\in\mathbb{R}^{|\mathcal{E}_{r}|\times 2}$. Each row in this matrix includes the free-flow travel time and the link capacity for a single link.
Furthermore, the original node features are often sparse and non-normalized. To address this issue, we employ a preprocessing step to encode the raw features into a lower-dimensional representation. This process not only captures the essential attributes of the data but also retains the semantic information. The generated node feature embedding size is $\bm{X}^{0}_{n}\in\mathbb{R}^{|\mathcal{V}|\times N_{v}}$, where $N_{v}$ is the embedding size. Similarly, the edge features are also normalized before being propagated in the message passing. Additionally, it should be noted that there is an overlap between the real links and virtual links in the heterogeneous graph because it is possible that a road and an OD demand both exist between the same node pair.
@@ 3.2 Graph Spatial Features Extraction
As discussed in Section 2.3, the key challenge in heterogeneous graphs lies in effectively aggregating diverse node and edge information. To address this challenge, we propose a novel sequential graph encoder for feature extraction and propagation. The sequential graph encoder is twofold: virtual graph encoder (V-Encoder) and real graph encoder (R-Encoder). Each part leverages attention mechanisms tailored to either virtual or real links, respectively. We will elaborate on these components in the following sections.
3.2.1 Virtual Graph Encoder (V-Encoder)
As the first step of spatial feature extraction, we employ a graph transformer-based attention mechanism on graph features to enhance the modeling capability. Specifically, we transform node features into key, query, and value matrices and then calculate the attention score between node pairs. Furthermore, since the virtual links are synthetically generated without inherent edge features, we introduce a learnable adaptive weight for virtual links, serving as their edge features. The adaptive weight for each node pair is derived by concatenating the two node features and passing through a position-wise feed-forward network (FFN). Mathematically, $L^{\mathrm{th}}$ V-Encoder can be expressed as:
$\bm{q}^{L,i},\bm{k}^{L,i},\bm{v}^{L,i}$ $=\bm{x}^{L}_{u}\ [\bm{W}_{q}^{L,i},\bm{W}_{k}^{L,i},\bm{W}_{v}^{L,i}],$ $\forall u\in\mathcal{V}$ (7)
$\beta_{e}^{L,i}$ $=\mathtt{FFN}\left([\bm{x}^{L,i}_{u}\oplus\bm{x}^{L,i}_{v}];\bm{W}^{L,i}_{\beta},\bm{b}^{L,i}_{\beta}\right),$ $\forall e=(u,v)\in\mathcal{E}_{v}$
$s_{e}^{L,i}$ $=\mathrm{exp}\left(\frac{\bm{q}_{u}^{L,i}\bm{k}_{v}^{L,i}}{\sqrt{d_{L}}}\beta_{e}^{L,i}\right),$ $\forall e=(u,v)\in\mathcal{E}_{v}$
where $\bm{x}^{L}_{u}$ is the feature embedding of node $u\in\mathcal{V}$ at $L^{\mathrm{th}}$ layer of V-Encoder. $\bm{q}_{(\cdot)}\in\mathbb{R}^{d_{L}}$, $\bm{k}_{(\cdot)}\in\mathbb{R}^{d_{L}}$, and $\bm{v}_{(\cdot)}\in\mathbb{R}^{d_{L}}$ is the query, key, and value vector at $i^{\mathrm{th}}$ head in V-Encoder where $d_{L}$ denotes the dimensionality of the feature vectors. $\bm{W}_{(\cdot)}$ and $\bm{b}_{(\cdot)}$ is the learnable parameters. $\beta_{e}^{L,i}\in\mathbb{R}$ and $s_{e}^{L,i}\in\mathbb{R}$ represents the learnable adaptive weight and the unnormalized attention score of the edge $e=(u,v)\in\mathcal{E}_{v}$ at $i^{\mathrm{th}}$ head of $L^{\mathrm{th}}$ layer, respectively. $\oplus$ is the concatenation operator. The introduction of edge-level adaptive weights is motivated by the fact that the observed variability in low-dimensional node embedding, in turn reflects the variability in node OD demand. Specifically, node pairs with higher OD demands should receive higher attention scores since they have more significant impacts on the flow distribution. By introducing the adaptive edge-level weight, the graph encoder can adjust the attention score among these node pairs and adaptively propagate the most relevant information through virtual links, thereby enriching the model’s contextual understanding. Then value vectors normalized by the attention scores, are processed through another position-wise feed-forward network, accompanied by layer normalization. Additionally, a residual connection supplements the final output of the layer, ensuring the integration of original input features with learned representations for enhanced model performance. Finally, the output of the V-encoder is obtained by concatenating the outputs from all attention heads:
$\bm{z}^{L,i}_{u}$ $=\sum_{v\in\mathcal{N}_{o}(u)}s_{(u,v)}^{L,i}\bm{v}_{v}^{L,i}/\sum_{v\in\mathcal{N}_{o}(u)}s_{(u,v)}^{L,i},$ (8)
$\bm{x}_{u}^{L+1,i}$ $=\bm{x}_{u}^{L,i}+\mathtt{LayerNorm}\left(\mathtt{FFN}\left(\bm{z}_{u}^{L,i};\bm{W}^{L,i}_{z},\bm{b}^{L,i}_{z}\right)\right),$ $\forall u\in\mathcal{V}$
$\bm{x}_{u}^{L+1}$ $=\left[\bm{x}_{u}^{L+1,0}\oplus\bm{x}_{u}^{L+1,1}\oplus\dots\oplus\bm{x}_{u}^{L+1,N_{h}}\right],$
where $\bm{z}^{L,i}_{u}$ represents the normalized weighted vector of node $u$ and $\mathcal{N}_{o}(u)$ represents all the outgoing nodes connected to $u$. $N_{h}$ represents the number of attention heads of the V-Encoder. To enhance the propagation of node features throughout the network, we sequentially stack multiple layers of the V-Encoder. The final output of this stacked V-Encoder, denoted as $H_{n}^{0}$, serves as the input for the subsequent encoder.
3.2.2 Real Graph Encoder (R-Encoder)
The R-Encoder is designed to enhance and complement the functionality of the V-Encoder. These two encoders share a similar architecture but slightly differ in the graph attention score mechanism. To illustrate, some nodes in the V-Encoder do not exchange messages with others due to the absence of virtual link connections. The R-Encoder addresses this issue by updating node features through real links, ensuring a comprehensive assessment of direct and indirect node relationships. From a mathematical perspective, the $M^{\mathrm{th}}$ layer of R-Encoder is expressed as follows:
$\bm{q}^{M,j},\bm{k}^{M,j},\bm{v}^{M,j}$ $=\bm{h}^{M}_{u}\ [\bm{W}_{q}^{M,j},\bm{W}_{k}^{M,j},\bm{W}_{v}^{M,j}],$ $\forall u\in\mathcal{V}$ (9)
$s_{e}^{M,j}$ $=\mathrm{exp}\left(\sum_{p=1}^{P}\frac{\bm{q}_{u}^{M,j}\bm{k}_{v}^{M,j}}{\sqrt{d_{L}}}\beta^{r}_{e,p}\right),$ $\forall e=(u,v)\in\mathcal{E}_{r}$
$\bm{z}^{M,j}_{u}$ $=\sum_{v\in\mathcal{N}_{o}(u)}s_{(u,v)}^{M,j}\bm{v}_{v}^{M,j}/\sum_{v\in\mathcal{N}_{o}(u)}s_{(u,v)}^{M,j},$ $\forall u\in\mathcal{V}$
$\bm{h}_{u}^{M+1,j}$ $=\bm{h}_{u}^{M,j}+\mathtt{LayerNorm}\left(\mathtt{FFN}\left(\bm{z}_{u}^{M,j};\bm{W}^{M,j}_{z},\bm{b}^{M,j}_{z}\right)\right),$ $\forall u\in\mathcal{V}$
$\bm{h}_{u}^{M+1}$ $=\left[\bm{h}_{u}^{M+1,0}\oplus\bm{h}_{u}^{M+1,1}\oplus\dots\oplus\bm{h}_{u}^{M+1,N_{h}}\right],$ $\forall u\in\mathcal{V}$
where $\bm{h}^{M}_{u}$ is the feature embedding of node $u\in\mathcal{V}$ at $M^{\mathrm{th}}$ layer of R-Encoder. $\bm{q}_{(\cdot)}\in\mathbb{R}^{d_{L}}$, $\bm{k}_{(\cdot)}\in\mathbb{R}^{d_{L}}$, and $\bm{v}_{(\cdot)}\in\mathbb{R}^{d_{L}}$ is the query, key, and value matrices at $j^{\mathrm{th}}$ head of R-Encoder. $\bm{W}_{(\cdot)}$ and $\bm{b}_{(\cdot)}$ are the learnable parameters. $\beta^{r}_{e,p}$ represents the $p^{\mathrm{th}}$ normalized edge feature of link $e\in\mathcal{E}_{r}$. Despite the subtle difference in the attention mechanism, the node features are propagated in two distinct patterns in the V-Encoder and R-Encoder. To illustrate, the V-Encoder captures the long-range dependency between nodes and integrates the contextual information from non-adjacent nodes. On the contrary, the R-Encoder captures the local topological relationships. Compared with the homogeneous graph with only real links, the additional feature propagation through virtual links can be considered as a dimension-reduction technique to reduce the number of hops required for distant nodes to gather messages. Consequently, it requires fewer GNN layers for effective feature aggregation and following edge prediction. Similar to the V-Encoder, the multiple R-Encoder layers are stacked together, and the output of the last R-Encoder $\bm{O}$, serves as the input for link flow prediction.